OpenStudy (anonymous):

What is the sum of the first seven terms of the geometric series 3 − 6 + 12 − . . . ?

rishavraj (rishavraj):

see its a Infinite Geometric Series..... a_1 = 3 r = -2

OpenStudy (anonymous):

Ya, im just trying to find the 7th term.

OpenStudy (anonymous):

so like s7=?

rishavraj (rishavraj):

in a geomtric series \[a_n = a_1 r^{n-1}\] and \[S_n =a_1\frac{ 1 - r^n }{ 1-r }\]

rishavraj (rishavraj):

so u want to find the sum of first 7 terms

OpenStudy (anonymous):

yes

rishavraj (rishavraj):

okkk it means \[S_7 = 3(\frac{ 1 - (-2)^7 }{ 1 - (-2) })\]

OpenStudy (anonymous):

oh alright let me plug that in

rishavraj (rishavraj):

go on nd solve S_7

OpenStudy (anonymous):

Ok thank you for your help!

rishavraj (rishavraj):

u welcome :)